# Enhancing OCR for Sino–Vietnamese Language Processing via Fine-Tuned PaddleOCRv5

### v1

@@ Abstract
Recognizing and processing Classical Chinese (Han–Nom) texts play a vital role in digitizing Vietnamese historical documents and enabling cross-lingual semantic research. However, existing OCR systems struggle with degraded scans, non-standard glyphs, and handwriting variations common in ancient sources such as Khâm Định Việt Sử Thông Giám Cương Mục. In this work, we propose a fine-tuning approach for PaddleOCRv5 to improve character recognition on Han–Nom texts. We retrain the text recognition module using a curated subset of ancient Vietnamese Chinese manuscripts, supported by a full training pipeline covering preprocessing, LMDB conversion, evaluation, and visualization. Experimental results show a significant improvement over the base model — exact accuracy rises from 37.5% to 50.0%, particularly under noisy image conditions. Furthermore, we develop an interactive demo that visually compares pre- and post-fine-tuning recognition results, facilitating downstream applications such as Han–Vietnamese semantic alignment, machine translation, and historical linguistics research. The demo is available at: https://huggingface.co/spaces/MinhDS/Fine-tuned-PaddleOCRv5
Keywords: OCR, PaddleOCR, Han–Nom, Classical Chinese, Vietnamese linguistics, fine-tuning, semantic processing.
@@ I. Introduction and Related Work
In the context of digital preservation and cultural heritage research, the construction of electronic corpora for linguistic studies, machine translation, and the analysis of ancient texts has become increasingly crucial. In Vietnam, a particularly valuable source of historical material is the body of Classical Chinese (Han) documents, exemplified by the monumental work Khâm Định Việt Sử Thông Giám Cương Mục [1], compiled during the Nguyễn dynasty with both Han and Vietnamese (Qu´ôc ngữ) versions. [^1] https://quangduc.com/a4672/kham-dinh-su-viet-thong-giam-cuong-muc-pdf. However, the automatic extraction of information from these historical sources remains technically challenging, particularly in the task of character recognition from degraded manuscript images.
Figure 1: Examples of OCR errors in ancient Chinese texts: noisy characters, missing lines, and sentence merging.
@@
Based on error analysis from our preliminary study (see Fig. 1), we observe that the default PP-OCRv5 model is not optimized for ancient Han–Nom data. This paper therefore proposes a fine-tuning approach for PaddleOCRv5 using a curated dataset of handwritten Han–Nom texts. Our goal is to enhance recognition performance for ancient Vietnamese Chinese documents, enabling applications in knowledge extraction, translation studies, and cross-lingual semantic research. The main contributions of this paper are as follows:
  • We construct a complete training pipeline for character recognition on ancient Han–Nom data, including image preprocessing, LMDB data conversion, and fine-tuning of the PP-OCRv5 model with optimized hyperparameters.
@@ 2.1 Fine-Tuning the Character Recognition Module of PP-OCRv5
In this study, we focus on fine-tuning the recognition module of PP-OCRv5 [4] (PP-OCRv5_rec) to improve the recognition performance for Han–Nom (Classical Chinese–Vietnamese) texts (see Fig. 2). The recognition module of PP-OCRv5 adopts a dual-branch architecture consisting of:
@@
$\mathcal{L}_{\text{total}}=\lambda_{1}\cdot\mathcal{L}_{\text{CTC}}+\lambda_{2}\cdot\mathcal{L}_{\text{KD}},$ (1)
@@ 2.2 Training Data and Preprocessing
Images used for fine-tuning 1,000,000
@@ 3.2 Main Results
Table 2: Recognition performance before and after fine-tuning.
@@
Figure 5: Recognition examples before (left) and after fine-tuning (right).
@@
Does fine-tuning improve character recognition on ancient Chinese texts? Yes. Fine-tuning yields a substantial gain in recognition performance, with exact accuracy increasing from 37.5% to 50.0%. This confirms that adapting the OCR model to domain-specific data (ancient Chinese texts) effectively enhances accuracy, even under limited data and suboptimal imaging conditions.
@@ IV. System Implementation
To demonstrate the practical effectiveness of the fine-tuned model, we developed an interactive demo system deployed on Hugging Face Spaces [^4] https://huggingface.co/spaces. The deployment uses a basic CPU environment (2 vCPUs, 16 GB RAM) and allows users to upload images of ancient Chinese text to obtain recognition results in real time.
Figure 6: Demo interface for recognizing ancient Chinese text using the fine-tuned model.
@@ V. Conclusion
Minh Hoang Nguyen was responsible for writing the Abstract, describing the model training and experimental setup (Training and Experiments), implementing and documenting the interactive demo system (System Demonstration), and contributing to the overall writing and editing of the paper. Su Nguyen Thiet contributed to the Introduction and Related Work sections, including the analysis of previous limitations and the description of the proposed PaddleOCR-based approach. He also led the construction of the training dataset, summarized the experimental outcomes, and jointly participated in the writing and final revision of the paper.

### v2

@@ Abstract
Recognizing and processing Sino-Nom (Han–Nom) texts play a vital role in digitizing Vietnamese historical documents and enabling cross-lingual semantic research. However, existing OCR systems struggle with degraded scans, non-standard glyphs, and handwriting variations common in ancient sources. In this work, we propose a fine-tuning approach for PaddleOCRv5 to improve character recognition on Han–Nom texts. We retrain the text recognition module using a curated subset of ancient Sino-Nom manuscripts, supported by a full training pipeline covering preprocessing, LMDB conversion, evaluation, and visualization. Experimental results show a significant improvement over the base model — exact accuracy rises from 37.5% to 50.0%, particularly under noisy image conditions. Furthermore, we develop an interactive demo that visually compares pre- and post-fine-tuning recognition results, facilitating downstream applications such as Han–Vietnamese semantic alignment, machine translation, and historical linguistics research. The demo is available at: https://huggingface.co/spaces/MinhDS/Fine-tuned-PaddleOCRv5
Keywords: OCR, PaddleOCR, Han–Nom, Sino-Nom, Vietnamese linguistics, fine-tuning, semantic processing.
@@ I. Introduction and Related Work
In the context of digital preservation and cultural heritage research, the construction of electronic corpora for linguistic studies, machine translation, and the analysis of ancient texts has become increasingly crucial. In Vietnam, a particularly valuable source of historical material is the body of Sino-Nom documents, with many characters borrowed or adapted from ancient Chinese texts, exemplified by the monumental work Khâm Định Việt Sử Thông Giám Cương Mục [1], compiled during the Nguyễn dynasty with both Han and Vietnamese (Qu´ôc ngữ) versions. [^1] https://quangduc.com/a4672/kham-dinh-su-viet-thong-giam-cuong-muc-pdf. However, the automatic extraction of information from these historical sources remains technically challenging, particularly in the task of character recognition from degraded manuscript images.
Figure 1: Examples of OCR errors in ancient Han–Nom texts: noisy characters, missing lines, and sentence merging.
Figure 2: An example of an alignment error caused by incorrect character mapping due to OCR mistakes.
@@
Based on error analysis from our preliminary study (see Fig. 1, and Fig. 2), we observe that the default PP-OCRv5 model is not optimized for ancient Han–Nom data. This paper therefore proposes a fine-tuning approach for PaddleOCRv5 using a curated dataset of handwritten Han–Nom texts. Our goal is to enhance recognition performance for Sino-Nom documents, enabling applications in knowledge extraction, translation studies, and cross-lingual semantic research. The main contributions of this paper are as follows:
  • We present a fine-tuned OCR system specifically for Han–Nom (Sino–Vietnamese) texts, leveraging a curated dataset of Vietnamese and broader Asian historical documents, addressing noise, degradation, and non-standard glyph variations common in ancient manuscripts.
@@
  • Experimental results confirm that our approach effectively improves OCR accuracy on Han–Nom datasets.
@@ 2.1 Fine-Tuning the Character Recognition Module of PP-OCRv5
In this study, we focus on fine-tuning the recognition module of PP-OCRv5 [4] (PP-OCRv5_rec) to improve the recognition performance for Han–Nom (Sino–Vietnamese characters) texts (see Fig. 3). The recognition module of PP-OCRv5 adopts a dual-branch architecture consisting of:
@@
$\mathcal{L}_{\text{total}}=\lambda_{1}\cdot\mathcal{L}_{\text{CTC}}+\lambda_{2}\cdot\mathcal{L}_{\text{KD}}$ (1)
@@ 2.2 Training Data and Preprocessing
Images used for fine-tuning 400,000
@@ 3.2 Main Results
Table 2: Performance results before and after fine-tuning.
@@
Figure 6: OCR Recognition examples before fine-tuning (left) and after fine-tuning (right).
@@
Does fine-tuning improve character recognition on ancient Chinese (Sino-Nom) texts? Yes. Fine-tuning yields a substantial gain in recognition performance, with exact accuracy increasing from 37.5% to 50.0%. This confirms that adapting the OCR model to domain-specific data (ancient Chinese texts) effectively enhances accuracy, even under limited data and suboptimal imaging conditions.
@@ IV. System Implementation
To demonstrate the practical effectiveness of the fine-tuned model, we developed an interactive demo system deployed on Hugging Face Spaces [^4] https://huggingface.co/spaces. The deployment uses a basic CPU environment (2 vCPUs, 16 GB RAM) and allows users to upload images of Sino-Nom text to obtain recognition results in real time.
Figure 7: Demo interface for recognizing ancient Sino-Nom text using the fine-tuned model.
@@ V. Conclusion
Minh Hoang Nguyen was Project Administration; Conceptualization; Writing – Original Draft; System Implementation; Methodology; Review & Editing. Su Nguyen Thiet contributed to Methodology; Data Curation; Writing – Introduction & Related Work; Experimental Summary; Writing – Review & Editing.
Figure 8: Construction of Building the bilingual Han–Vietnamese pipeline.
Appendix A Appendix A: Alignment Pipeline and OCR Fine-Tuning Impact
In this appendix, we provide a detailed description of the Han–Vietnamese sentence alignment process and demonstrate how fine-tuning the PaddleOCRv5 model improves alignment quality. This section complements the main paper by highlighting technical details of preprocessing, alignment, and empirical results. For our experiments, we use the exemplary historical work Khâm Định Việt Sử Thông Giám Cương Mục [1], based on PDF scans of Sino-Nom texts:
  • Original Han text: Stored at the National Library of Vietnam, in PDF image format, carved in 1884.
  • Vietnamese translation [^6] https://quangduc.com/a4672/kham-dinh-su-viet-thong-giam-cuong-muc-pdf: Provided by the Vietnam Institute of Historical Studies. Official translation, Education Publishing House, Hanoi, 1998.
Appendix B A.1 Alignment Pipeline
The construction of the bilingual Han–Vietnamese corpus follows three main stages (see Figure 8):
  1. 1.
    Han text processing: PDF scans of 51 historical volumes are converted into page images, preprocessed (resizing, denoising, binarization), and processed with OCR. Post-processing involves box correction, line merging, and sequence ordering according to traditional vertical writing (right-to-left, top-to-bottom). The result is a cleaned sequence of Han sentences $\mathcal{H}=\{h_{1},h_{2},\dots,h_{N}\}$.
  2. 2.
    Vietnamese translation processing: PDF translations are extracted and cleaned to remove noise (titles, footnotes, page numbers) and split into sentences $\mathcal{V}=\{v_{1},v_{2},\dots,v_{M}\}$, ensuring consistent structure suitable for alignment.
  3. 3.
    Sentence alignment: We apply Bertalign [6] to map Han sentences to Vietnamese sentences. Bertalign computes semantic embeddings for each sentence and uses dynamic programming to find the optimal 1–1, 1–n, or n–1 alignment:
    $\text{alignment}(\mathcal{H},\mathcal{V})=\arg\min\sum c(x,y)$
    where $c(x,y)$ is the cost function based on cosine similarity between embedding vectors, adjusted for sentence count in multi-sentence segments. The output is a set of aligned sentence pairs $\mathcal{A}=\{(h_{i},v_{j})\}$ with confidence scores $s_{ij}\in[0,1]$.
Appendix C A.2 Integration of OCR Fine-Tuning
OCR errors in Han text significantly affect alignment quality. Misrecognized characters can lead to incorrect semantic embeddings, resulting in lower alignment confidence or mismatched sentence pairs. To address this, we fine-tuned the character recognition module of PaddleOCRv5(see Section II in main paper).
After fine-tuning, the OCR output exhibits higher character accuracy and confidence, which directly improves downstream alignment. Specifically:
  • Error reduction: Character misrecognition in low-quality scanned pages decreases from 62.5% to 50%.
  • Alignment robustness: More sentences are correctly matched to their Vietnamese counterparts, particularly in cases involving rare or archaic characters.
Appendix D A.3 Experimental Results
We conducted experiments comparing alignment performance using OCR outputs from the baseline and fine-tuned models. Table 3 summarizes the findings. Metrics include:
  • Aligned sentence ratio: Proportion of Han sentences successfully mapped to Vietnamese.
  • Average confidence: Mean Bertalign alignment score across all sentence pairs.
  • Exact 1–1 alignment rate: Fraction of sentences aligned in a strict 1–1 manner.
Table 3: Effect of OCR fine-tuning on Han–Vietnamese sentence alignment.
Metric
Baseline
OCR
Fine-Tuned
OCR
Aligned sentence ratio 61.8% 67.6% ↑
Average alignment confidence 0.783 0.842 ↑
Exact 1–1 alignment rate 46.2% 50.4% ↑
Total aligned pairs 15,987 17,455 ↑
The results indicate a clear improvement in both the quantity and quality of aligned sentence pairs after fine-tuning OCR. Notably:
  • More sentences with rare or complex characters are correctly aligned.
  • Average semantic similarity scores increase, reflecting higher alignment confidence.
  • Downstream applications, such as machine translation or bilingual corpus construction, benefit from more reliable sentence pairs.
Fine-tuning the OCR model on Chinese Document Text Recognition Dataset substantially reduces recognition errors, which translates into improved semantic embeddings for alignment. Consequently, the Bertalign-based pipeline yields a larger number of high-confidence aligned sentences. This appendix demonstrates the direct impact of OCR enhancement on bilingual corpus quality, confirming that preprocessing and OCR fine-tuning are crucial steps in building reliable Han–Vietnamese datasets.